# "I understand why I got this grade": Automatic Short Answer Grading with Feedback

### v1

@@ Abstract
The demand for efficient and accurate assessment methods has intensified as education systems transition to digital platforms. Providing feedback is essential in educational settings and goes beyond simply conveying marks as it justifies the assigned marks. In this context, we present a significant advancement in automated grading by introducing Engineering Short Answer Feedback (EngSAF)—a dataset of student answers accompanied by reference answers and questions for the Automatic Short Answer Grading (ASAG) task. The EngSAF dataset is meticulously curated to cover a diverse range of subjects, questions, and answer patterns from multiple engineering domains. We leverage state-of-the-art large language models’ (LLMs) generative capabilities with our Label-Aware Synthetic Feedback Generation (LASFG) strategy to include feedback in our dataset. This paper underscores the importance of enhanced feedback in practical educational settings, outlines dataset annotation and feedback generation processes, conducts a thorough EngSAF analysis, and provides different LLMs-based zero-shot and finetuned baselines for future comparison111Our code and dataset are available at https://github.com/dishankaggarwal/EngSAF. Additionally, we demonstrate the efficiency and effectiveness of the ASAG system through its deployment in a real-world end-semester exam at the Indian Institute of Technology Bombay (IITB), showcasing its practical viability and potential for broader implementation in educational institutions.
1 Introduction
Technology integration in education has resulted in transformative changes, redefining traditional pedagogical approaches and assessment methodologies. Effective education relies on feedback and explanations provided during assessments to ensure quality learning outcomes Shute (2008). Grading questions in tests and examinations have proven to be a good measure to assess student’s learning and understanding of a topic or a subject. An exam could include various question types, such as multiple choice, fill-in-the-blank, short answers, essays, etc. Among these question types, short answers and essays are more complicated to analyze than multiple-choice or fill-in-the-blank type questions due to flexibility and natural language in the response. Automating the grading process becomes crucial, especially in countries with extremely high student-to-teacher ratios, as it can significantly reduce instructor’s workloads and improve the assessment process. Significant advancements have been achieved in this field in recent years, primarily after the introduction of SemEval Semantic Textual Similarity (STS) task Agirre et al. (2012, 2013, 2014, 2015).
@@ 1 Introduction
This challenge can be approached as a machine learning issue, where the objective is to grade a student’s response based on how similar it is to the reference answers. However, simply assigning a score or label to a learner’s response is often inadequate in practical educational contexts. Ahea et al. (2016) highlights the value and effectiveness of feedback in improving students’ learning and professionalizing teaching in higher education.
1.1 Problem Statement
@@
Given a question, a reference answer, and a student’s answer, the aim is to provide content-focused elaborated feedback and assign a label indicating the degree of correctness in the student’s answer compared to the reference answer. Here, we focus on questions where the answer type is a sentence or a short paragraph. This task involves evaluating the alignment between the student’s and reference answers. Figure 1 illustrates the problem statement, showing the input and output using an example from the EngSAF dataset.
@@
1.2 Motivation
The increasing demand for technology in education has led to a need for more efficient and effective methods of grading short-answer assessments. In the case of short answer grading, feedback with clear explanations goes beyond simply conveying a grade. It offers valuable insights into student’s strengths, weaknesses, and areas of improvement. However, the effectiveness of feedback is largely unexplored due to the lack of public, content-centered, elaborated feedback datasets in different domains. These datasets are crucial for training and developing automated feedback systems that can provide personalized and nuanced feedback. The need for a more efficient and effective ASAG system incorporating feedback alongside grade has arisen due to these limitations, and this is where the Engineering Short Answer Feedback (EngSAF) dataset steps in, consisting of questions students answer from multiple engineering domains for the ASAG task.
Our Contributions are:
1.
EngSAF dataset containing around 5.8K student responses to 119 questions from multiple engineering domains along with synthetically generated feedback explaining the assigned grade for the task of ASAG. To the best of our knowledge, this is the first dataset containing questions and responses from multiple engineering domains. (Section 3)
2.
Benchmark scores on the EngSAF dataset using different Large Language Models (LLM) for future comparison and research. (Section 5)
3.
Real-world deployment of the EngSAF fine-tuned ASAG model in an end-semester exam at Indian Institute of Technology, Bombay (IITB) (Section 6).
2 Related work
2.1 Automatic Short Answer Grading (ASAG)
ASAG is an essential area of research that has garnered significant attention in recent years. Several approaches have been proposed for traditional ASAG, ranging from rule-based methods to more sophisticated machine-learning techniques. One early approach for ASAG was based on keyword or pattern matching, where the presence or absence of certain keywords in the student’s answer was used to determine its accuracy. Mitchell et al. (2002); Sukkarieh et al. (2004); Nielsen et al. (2009).To overcome these limitations, researchers have developed more sophisticated methods that use natural language processing (NLP) techniques. One such method is based on Latent Semantic Analysis (LSA), which represents texts as high-dimensional vectors and compares them to the reference answers using cosine similarity LaVoie et al. (2020). In a related study, the task of ASAG is addressed by incorporating features such as answer length, grammatical correctness, and semantic similarity in comparison to reference answers Sultan et al. (2016).
More recently, deep learning models such as convolutional neural networks (CNNs) and recurrent neural networks (RNNs) have been applied to ASAG’s task. These models are trained on large amounts of annotated data and can capture the semantic relationships between words in a student’s answer, and the reference answers Surya et al. (2019); Zhang et al. (2022). Pre-trained Language Models (PLMs), such as BERT Devlin et al. (2018), GPT Radford et al. (2019), RoBERTa Liu et al. (2019), DistillBERT Sanh et al. (2019), and ALBERT Lan et al. (2019) have performed exceptionally well for various tasks in NLP, also used for the task of ASAG. Sentence-BERT or SBERT Reimers and Gurevych (2019) performs exceptionally well on current publicly available traditional ASAG datasets Condor et al. (2021). A common limitation of the prior work is that only grades are assigned, and no feedback is provided. We seek to fill this gap with the introduction of the EngSAF dataset.
2.2 Short Answer Grading Datasets
Several publicly accessible datasets have been curated, each designed to facilitate research and benchmarking in the ASAG task. In 2009, Mohler and Milhalcea published a short answer dataset on a data structures course that contained 630 records Mohler and Mihalcea (2009). In 2011, they published an extended dataset on the same course that contained 2273 records Mohler et al. (2011), named the Texas Extended dataset. To complete the assignment of student-authored answer annotation, the Beetle corpus is being created Nielsen et al. (2008). Another dataset that is publicly available on Kaggle and could be used for the task of ASAG is provided by Hewlett Foundation named ASAP-AES222https://www.kaggle.com/c/asap-aes/data (Automated Assessment Prize Competition for Essay Scoring).
However, all aforementioned datasets are conventional ASAG datasets solely consisting of grades or scores. Our extensive literature survey includes only one short answer grading dataset with content-focused elaborated feedback, i.e., Filighera et al. (2022), which introduces an inaugural dataset for short-answer feedback comprising bilingual responses in English and German. Human annotations were meticulously collected and refined to uphold the feedback’s quality. However, this dataset is constrained by its limited number of student responses () in English and its exclusive focus on questions from a single domain, specifically computer network queries, thereby lacking diversity across different engineering domains. To address these limitations, we introduced EngSAF, which contains data points and questions from multiple engineering domains.
3 Engineering Short Answer Feedback (EngSAF) Dataset
This dataset contains 119 questions drawn from different undergraduate and graduate engineering courses, accompanied by approximately 5.8k student responses. An Instructor provided correct answer/ reference answer also accompanies each question. These questions and answers have been taken from actual quizzes/ exams from the Indian Institute of Technology, Bombay (IITB), covering a diverse range of 25 courses. The questions and responses span across multiple domains, including image processing, water quality management, and operating systems, to name a few. Technical Appendix contains the complete list of courses included in the dataset. Based on the instructor’s assigned marks, each response has been categorized as “correct,” “partially correct,” or “incorrect.”
@@ 3 Engineering Short Answer Feedback (EngSAF) Dataset
Table 1 provides an example from the EngSAF dataset showcasing a question, its corresponding reference answer, and three different student responses alongside their associated output labels and synthetically generated feedback/Explanation. Technical Appendix contains more examples from the EngSAF dataset. The dataset is applicable for both traditional automatic short answer grading and the generation of elaborated feedback.
@@ 3.1 Dataset Construction
EngSAF contains questions and student responses from different undergraduate and graduate engineering courses. The instruction provided a correct answer/ reference answer accompanying each question. The Instructor/ Teaching Assistant (TA) assigns marks to each student’s answer from their respective course. The students’ responses and grades are obtained from a reputed engineering university. The quality assessment of the proposed dataset is covered in subsection 3.5. If the maximum mark for a question is and a student’s answer is graded marks by the instructor, then the output label is annotated according to the scheme in Table 2.
@@
3.2 Challenges and Requirements
Diversity in student responses: Human language is inherently subjective, and different individuals may interpret the same content differently. This subjectivity extends to the grading process, making it challenging to provide universally applicable feedback.
Response Variability: Short answer responses can vary significantly regarding quality, coherence, and relevance. Some responses may be well-structured and articulate, while others may be incomplete or contain grammatical errors. Designing feedback that caters to this variability in response quality poses a considerable challenge.
Feedback Impact: Wrong feedback can negatively shape learners’ perceptions of themselves and their abilities. If the feedback consistently highlights their shortcomings or mistakes, they may internalize these negative perceptions and develop a fixed mindset about their capabilities. This can hinder their willingness to take risks, seek challenges, and persist in facing difficulties.
@@ 3.3 Label Aware Synthetic Feedback Generation (LASFG)
Leveraging the advanced language generative and reasoning capabilities of state-of-the-art Large Language Models (LLMs) like Gemini333https://gemini.google.com/ and ChatGPT, we enhance educational assessment with feedback, particularly in Automatic Short Answer Grading (ASAG). The approach involves utilizing Gemini’s ability to comprehend input prompts consisting of a question, a student’s answer, a reference answer, and the corresponding grading label provided by an instructor or teaching assistant to generate content-focused elaborated feedback. The generated feedback covers the reasoning or explanation of the Gold output label. Grammarly444https://www.grammarly.com/ then checks the synthetically generated feedback to remove any grammatical errors. Prompt details used for synthetic feedback generation can be found in the Technical Appendix 10.1. The quality estimation of the generated feedback is shown in Section 3.5
@@ 3.4 Corpus Statistics
Following Dzikovska et al. (2013), we partitioned the data into training sets, comprising 70% of the dataset, as well as unseen answers (16%) and unseen questions (14%) for test sets as shown in table 3. The test split with unseen answers includes fresh responses to the questions used in training. In contrast, the test split with unseen questions comprises entirely new questions for testing the model’s ability to generalize to new questions without prior exposure. Technical Appendix 10.3 contains insight about insights about the distribution of text length for student answers, reference answers, questions, and feedback in the EngSAF dataset.
@@ 3.5 Quality Estimation
To show the reliability and credibility of our dataset and synthetically generated feedback, we randomly sampled 300 data points and equally distributed them across the output label. This sampled data is a good representation of the EngSAF dataset.
Each synthetically generated feedback was scored by three human annotators over three aspects. Each aspect is scored on a scale (1-5), with a high score indicating a better response. Each feedback is analyzed on the following aspects.
@@
Fluency and Grammatical Correctness: This aspect tests whether the generated feedback is fluent in English and grammatically correct.
@@
Feedback Correctness/Accuracy: This aspect assesses the overall quality of the generated feedback regarding content, relevance, quality, and explanation for the assigned grade.
@@
Emotional Impact: This aspect assesses how feedback affects the learner’s emotional state. Annotators are tasked with rating whether the feedback avoids triggering negative emotions or impacts by refraining from using words such as "fail" that may evoke feelings of discouragement or distress in the learner.
@@
| Fluency & Grammatical Correctness | 4.73 |
| Feedback Correctness/Accuracy | 4.55 |
| Emotional Impact | 4.61 |
Three human annotators evaluate the correctness of each output label for sampled data points, achieving an accuracy of 98% and pair-wise average Cohen’s Kappa () score of 0.65 (substantial agreement), showcasing the high reliability of the assigned output label. Table 4 shows the average score over all the designed aspects to measure the reliability of the proposed dataset. Human evaluation across various designed aspects consistently yields an average score greater than 4.5 out of 5, which underscores the reliability and quality of the EngSAF dataset. The details of the human annotators can be found in Technical Appendix.

### v2

# “I understand why I got this grade”: Automatic Short Answer Grading (ASAG) with Feedback

@@ Abstract
In recent years, there has been a growing interest in using Artificial Intelligence (AI) to automate student assessment in education. Among different types of assessments, summative assessments play a crucial role in evaluating a student’s understanding level of a course. Such examinations often involve short-answer questions. However, grading these responses and providing meaningful feedback manually at scale is both time-consuming and labor-intensive. Feedback is particularly important, as it helps students recognize their strengths and areas for improvement. Despite the importance of this task, there is a significant lack of publicly available datasets that support automatic short-answer grading with feedback generation. To address this gap, we introduce Engineering Short Answer Feedback (EngSAF), a dataset designed for automatic short-answer grading with feedback. The dataset covers a diverse range of subjects, questions, and answer patterns from multiple engineering domains and contains data points. We incorporate feedback into our dataset by leveraging the generative capabilities of state-of-the-art large language models (LLMs) using our Label-Aware Synthetic Feedback Generation (LASFG) strategy. This paper underscores the importance of enhanced feedback in practical educational settings, outlines dataset annotation and feedback generation processes, conducts a thorough EngSAF analysis, and provides different LLMs-based zero-shot and finetuned baselines for future comparison111Dataset and code are open-sourced at: https://github.com/dishankaggarwal/EngSAF. The best-performing model (Mistral-7B) achieves an overall accuracy of 75.4% and 58.7% on unseen answers and unseen question test sets, respectively. Additionally, we demonstrate the efficiency and effectiveness of our ASAG system through its deployment in a real-world end-semester exam at a reputed institute. There, we achieve an output label accuracy of 92.5% along with feedback correctness and emotional impact scores above 4.5 (out of 5) on human evaluation, thus showcasing its practical viability and potential for broader implementation in educational institutions.
Keywords:
LLM Automatic Short Answer Grading EngSAF Feedback1 Introduction
Technology integration in education, particularly artificial intelligence (AI), has resulted in transformative changes that have redefined traditional pedagogical approaches and assessment methodologies. Effective education relies on feedback and explanations provided during assessments to ensure quality learning outcomes [30]. In the context of education, educators use different kinds of assessments, including formative, summative, and diagnostic222https://www.niu.edu/citl/resources/guides/instructional-guide/formative-and-summative-assessment.shtml, each of which serves a vital role in a student’s learning journey. This work deals with summative assessments, which are essential in determining whether the student has achieved the learning goals of a course once it is over. Grading summative assessments often involve grading short answers and essays, which are more complicated in nature due to the flexibility and natural language in the response. Hence, automating the grading process for such questions becomes crucial, especially in countries with extremely high student-to-teacher ratios, as it can significantly reduce instructor’s workloads and improve the assessment process. For instance, during the COVID-19 pandemic, educators faced increased stress and exhaustion due to overloaded grading burden [15, 25].
@@
This challenge can be approached as a machine learning problem, where the objective is to grade a student’s response based on how similar it is to the reference answers. However, simply assigning a score or label to a learner’s response is often inadequate in practical educational contexts [13, 31]. Ahea et al. [5] highlights the value and effectiveness of feedback in improving student’s learning and standardizing teaching in higher education. Although earlier work formulated this problem as the Semantic Textual Similarity (STS) task [3, 4, 2, 1], recent works acknowledge the importance of feedback generation along with the grading task [12]. Hence, together with the grading task and the feedback generation task, researchers have formulated the problem of automatic short answer grading (ASAG) with feedback.
1.1 Automatic Short Answer Grading (ASAG) with Feedback
@@
The ASAG with feedback problem is as follows: given a question, a reference answer, and a student’s answer, the aim is to assign a label indicating the degree of correctness in the student’s answer compared to the reference answer and provide content-focused elaborated feedback/explanation for the same. Note that the degree of correctness is limited to only three labels here, namely correct, partially correct, and incorrect. In this problem, we only focus on questions where the answers are short, varying between a sentence and a short paragraph. This task involves evaluating the alignment between the student’s answer and the reference answers which is expressed via the degree of correctness along with proper reasoning as to why that label is assigned.
Figure 1 illustrates the problem statement, showing the input and output using an example from the EngSAF dataset.
1.2 Motivation
With recent advancements in state-of-the-art (SOTA) Natural Language Processing (NLP) techniques, there is an increasing interest in applying such methodologies in the field of education, particularly in short-answer evaluations. The benefits of designing such AI-assisted systems that solve the ASAG with feedback problem are two-fold:
1.
The learners333We often use the terms student and learner synonymously. Also, we do the same for the terms graders, teachers, and educators. benefit by understanding their strengths and shortcomings from the content-elaborated feedback provided by the system.
2.
The educators benefit by having prior feedback on the student’s degree of correctness, along with a proper explanation for the same. This can reduce an educator’s cognitive load while grading such short answer questions.
However, researchers face challenges when designing such systems due to the lack of public, content-centered, elaborated feedback datasets in different domains. These datasets are crucial for training and developing automated grading systems that can provide feedback. Thus, this calls for a more efficient and effective dataset spanning multiple engineering domains along with proper grades and content-elaborated feedback. This is where we introduce the Engineering Short Answer Feedback (EngSAF) dataset.
Our contributions are:
1.
A Label Aware Synthetic Feedback Generation (LASFG) strategy to convert conventional automatic short answer grading datasets to the ones containing feedback. (Section 3.2)
2.
EngSAF dataset containing around 5.8K student responses to 119 questions from multiple engineering domains taken from real-life examinations at a reputed institute along with synthetically generated feedback using LASFG strategy explaining the assigned grade for the task of ASAG. To the best of our knowledge, this is the first dataset containing questions and responses from multiple engineering domains. The overarching goal of this work is to provide helpful feedback to students automatically. (Section 3)
3.
Benchmark scores on the EngSAF dataset using different Large Language Models (LLMs) for future comparison and research. (Section 5)
4.
Real-world deployment of the best-performing fine-tuned ASAG model in an end-semester exam at a reputed institute. (Section 6)
2 Related work
This section provides a comprehensive literature review, beginning with previous research on the ASAG problem, followed by an in-depth survey of existing short-answer grading datasets.
2.1 Automatic Short Answer Grading (ASAG)
ASAG is an essential area of research that has garnered significant attention in recent years. Several approaches have been proposed for traditional ASAG techniques, ranging from rule-based methods to more sophisticated machine-learning techniques. One early approach for solving the ASAG problem was based on keyword or pattern matching, where the presence or absence of certain keywords in the student’s answer was used to determine its accuracy [20, 32, 23]. To overcome these limitations, researchers have developed more sophisticated methods that use NLP-based techniques. One such method is based on Latent Semantic Analysis (LSA), which represents texts as high-dimensional vectors and compares them to the reference answers using cosine similarity [17]. In a related study, the task of ASAG is addressed by incorporating features such as answer length, grammatical correctness, and semantic similarity in comparison to reference answers [33].
More recently, deep learning models such as convolutional neural networks (CNNs) and recurrent neural networks (RNNs) have been applied to ASAG’s task. These models are trained on large amounts of annotated data and can capture the semantic relationships between words in a student’s answer and the reference answers [34, 36]. Pre-trained Language Models (PLMs), such as BERT [8], GPT [27], RoBERTa [18], DistillBERT [29], ALBERT [16], and SBERT[28] have performed exceptionally well for various NLP tasks, are widely used even for the task of ASAG on traditional ASAG datasets [7]. A more recent work by Fateen et al. [11] extends this problem by generating feedback while solving the ASAG problem using a RAG-based approach. However, a common limitation of all the prior works is that only grades are assigned, and none of the datasets provide a ground truth for the problem content-elaborated feedback generation.
2.2 Short Answer Grading Datasets
Again, when it comes to short answer grading datasets, researchers have carefully curated several publicly accessible datasets to facilitate research and benchmarking of different solutions on the ASAG task. In 2009, Mohler et al. [22] published a short answer dataset on a data structures course that contained 630 records. In 2011, the same authors [21] published an extended dataset on the same course that contained 2273 records and named it the Texas Extended dataset. Apart from this, Nielsen et al. [24] created the Beetle corpus, which contains fine-grained annotations of entailment relationships in the context of student answers to science assessment questions. Another dataset that is publicly available on Kaggle and could be used for the task of ASAG is provided by Hewlett Foundation named ASAP-AES444https://www.kaggle.com/c/asap-aes/data (Automated Assessment Prize Competition for Essay Scoring).
However, as mentioned above, all datasets are conventional ASAG datasets that solely consist of grades or scores. Our extensive literature survey includes only one short-answer grading dataset with content-focused elaborated feedback, i.e., [12], which introduces an inaugural dataset for this problem and comprises bilingual responses in English and German. Human annotations were meticulously collected and refined to uphold the quality of the feedback. However, this dataset is constrained by its limited number of student responses () in English and its exclusive focus on questions from a single domain, specifically computer network queries, thereby lacking diversity across different engineering domains. To address these limitations, we curate the EngSAF dataset, which contains data points and questions from multiple engineering domains.
3 Engineering Short Answer Feedback (EngSAF) Dataset
This dataset contains 119 questions drawn from different undergraduate and graduate engineering courses, accompanied by approximately 5.8k student responses. An instructor provided correct/reference answer also accompanies each question. These questions and answers have been taken from real-life examinations at a reputed institute, thus covering a diverse range of 25 courses. The questions and responses span across multiple domains, including image processing, water quality management, and operating systems, to name a few. The dataset has been curated using a well-known e-learning platform that has been extensively used and developed at a reputed institute. Based on the instructor’s assigned marks, each response has been categorized as “correct,” “partially correct,” or “incorrect.”
@@ 3 Engineering Short Answer Feedback (EngSAF) Dataset
Table 1 provides examples from the EngSAF dataset showcasing a question, its corresponding reference answer, and three different student responses alongside their associated output labels and synthetically generated feedback. The dataset is applicable for both traditional automatic short answer grading and the generation of elaborated feedback.
@@ 3.1 Dataset Construction
The EngSAF dataset contains questions and student responses from different undergraduate and graduate engineering courses from a reputed institute. The data was collected using a well-known e-learning platform that has been developed and is extensively used at this institute. Using the e-learning platform, the instructor provided a correct answer/ reference answer to each question. Additionally, the instructors and teaching assistants (TAs) have assigned marks to each student’s answer from their respective courses. The student’s responses and grades were then obtained from the e-learning platform. The quality assessment of the proposed dataset is covered in subsection 3.4.
@@
Each data point includes a question, student answer, and reference answer, along with a label indicating the level of correctness of the student’s response and detailed, content-specific feedback. If the maximum mark for a question is and a student’s answer is graded marks by the instructor, then the output label is annotated according to the scheme in Table 2.
@@ 3.2 Label Aware Synthetic Feedback Generation (LASFG)
Apart from the question, the student’s answer, and the reference answer, each data point contains a content-elaborated feedback part. To obtain this, we leverage the advanced language generative and reasoning capabilities of SOTA LLMs like Gemini555https://gemini.google.com/ and ChatGPT. The approach involves utilizing Gemini’s ability to comprehend input prompts consisting of a question, a student’s answer, a reference answer, and the corresponding grading label provided by an instructor or teaching assistant to generate content-focused elaborated feedback. The generated feedback covers the reasoning or explanation of the Gold output label. Based on initial observations of the generated feedback quality, we use gemini-1.5-flash model for feedback generation. To ensure the quality of the generated feedback, we performed quality estimation as described in Section 3.4.
@@ 3.3 Corpus Statistics
Following Dzikovska et al.’s approach [10], we partitioned the data into training sets, comprising 70% of the dataset, as well as unseen answers (16%) and unseen questions (14%) for test sets as shown in Table 3. The test split with unseen answers (UA) includes fresh responses to the questions used in training. In contrast, the test split with unseen questions (UQ) comprises entirely new questions for testing the model’s ability to generalize to new questions without prior exposure. The average text length (in tokens) for different fields in our dataset is as follows: Questions have an average length of 17.50 tokens, student answers average 25.85 tokens, reference answers are 26.47 tokens long on average, and feedback is the longest, with an average of 41.95 tokens.
@@ 3.4 Quality Estimation
To show the reliability and credibility of our dataset and synthetically generated feedback, we randomly sampled 300 data points and equally distributed them across the output label. This sampled data is a good representation of the EngSAF dataset. Each synthetically generated feedback was evaluated on three aspects by three proficient English-speaking annotators, all Master’s students. They were provided with detailed annotation guidelines and examples to ensure consistency.
Each aspect is scored on a scale (1-5), with a high score indicating a better response. Each feedback is analyzed on the following aspects.
@@
Fluency and Grammatical Correctness (FGC): This aspect checks if the feedback is fluent and grammatically correct in English.
@@
Feedback Correctness/Accuracy (FC): This aspect assesses the overall quality of the generated feedback regarding content, relevance, quality, and explanation for the assigned grade.
@@
Emotional Impact (EI): This aspect assesses how feedback affects the learner’s emotional state. Annotators are tasked with rating whether the feedback avoids triggering negative emotions or impacts by refraining from using words such as “fail” that may evoke feelings of discouragement or distress in the learner.
@@
| Fluency & Grammatical Correctness (FGC) | 4.73 |
| Feedback Correctness/Accuracy (FC) | 4.55 |
| Emotional Impact (EI) | 4.61 |
Three human annotators evaluate the correctness of each output label for sampled data points, achieving an accuracy of 98% and pair-wise average Cohen’s Kappa () score of 0.65 (substantial agreement), showcasing the high reliability of the assigned output label. Table 4 shows the average score over all the designed aspects to measure the reliability of the proposed dataset. Human evaluation across various designed aspects consistently yields an average score greater than 4.5 out of 5, which underscores the reliability and quality of the EngSAF dataset.
As part of our ablation studies, we have also performed an LLM-based evaluation by following the LLM-as-a-judge paradigm [38]. We conducted such an evaluation using the Gemini model to assess the quality of the generated feedback. Specifically, we instructed Gemini to evaluate feedback quality on the EngSAF test set, and we obtained average scores of 4.23 for feedback quality on unseen answers (UA) and 3.91 on unseen questions (UQ) out of 5. These scores reflect the model’s assessment of the feedback quality and provide an additional layer of validation to the reliability and quality of the dataset.